OpenStudy (anonymous):

(5x - 1)/2 + (-3x - 1)/4 = 9/2

OpenStudy (anonymous):

lets clear the fractions... multiply the whole equation by 4... \[\large \frac{5x-1}{2}+\frac{-3x-1}{4}=\frac{9}{2} \] \[\large 4\times(\frac{5x-1}{2}+\frac{-3x-1}{4}=\frac{9}{2}) \] \[\large 2(5x-1)+(-3x-1)=2*9 \] can you take it from here?

OpenStudy (anonymous):

im lost.. :/

OpenStudy (maheshmeghwal9):

use this property\[\LARGE{\color{red}{2(a+b)=2a+2b.}}\] This is what that @dpaInc has done:)

OpenStudy (anonymous):

10x - 1 + -6x-2 = 18 ? idk im confused

OpenStudy (maheshmeghwal9):

It means distributive law of property. Only difference is that @dpaInc has use 4 instead of 2

OpenStudy (maheshmeghwal9):

k! \[\color{green}{4x-3=18\implies x=\frac{21}{4}.}\]

OpenStudy (maheshmeghwal9):

this is the further solution:)

OpenStudy (anonymous):

so thats my answer? im so lost! haha i hate math

OpenStudy (maheshmeghwal9):

ya its ur answer :-)

OpenStudy (maheshmeghwal9):

sorry wait a minute u have multiplied the equation wrong :/

OpenStudy (maheshmeghwal9):

10x - 1 + -6x-2 = 18 is wrong. The right is 10x-2-3x-1=18. now do it.

OpenStudy (anonymous):

how would i solve that? :/

OpenStudy (anonymous):

7x-1= 18?

OpenStudy (maheshmeghwal9):

just as i did before for wrong eq.

OpenStudy (maheshmeghwal9):

no it is nt this try again

OpenStudy (anonymous):

idk how to do it then :/

OpenStudy (maheshmeghwal9):

-1-2=-3 so ur eq. will become 7x-3=18. x=21/7. & for ur information it is ur answer:-)